40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 8 11:22:35 2014, steve, Summary, SUS, oplev laser summary 
    Reply  Mon Aug 11 10:23:39 2014, steve, Summary, SUS, oplev laser summary updated 
       Reply  Thu Sep 18 10:08:07 2014, steve, Summary, SUS, oplev laser summary updated 
          Reply  Thu Mar 19 17:11:32 2015, steve, Summary, SUS, oplev laser summary updated 
          Reply  Fri Nov 20 09:45:39 2015, steve, Summary, SUS, oplev laser summary  oplev_summery.png
             Reply  Mon May 23 10:55:49 2016, steve, Summary, SUS, ITMX oplev laser replaced ITMXoplev.png
          Reply  Mon May 23 15:51:32 2016, steve, Summary, SUS, oplev laser summary updated oplSum.png
             Reply  Tue May 24 10:21:36 2016, ericq, Summary, SUS, ITMX Oplev loops 
             Reply  Wed Jul 27 11:40:48 2016, steve, Summary, SUS, oplev laser summary updated 
                Reply  Wed Jan 25 10:46:58 2017, steve, Summary, SUS, oplev laser summary updated oplev_sums.png
                   Reply  Wed Feb 8 10:16:39 2017, steve, Update, SUS, clipping ITMX oplev ITMX_oplev_clipping.jpgITMX_clipping.jpg
                      Reply  Fri May 18 07:41:01 2018, steve, Update, SUS, clipping ITMX oplev 
                   Reply  Mon Apr 10 09:58:35 2017, Steve, Summary, SUS, oplev laser summary updated 
                      Reply  Wed Jul 19 12:47:43 2017, Steve, Summary, SUS, oplev laser summary updated 
                         Reply  Wed Jul 19 18:24:15 2017, rana, Summary, SUS, oplev laser summary updated 
                            Reply  Fri Jul 21 08:05:07 2017, Steve, Summary, SUS, red, blue, green laser diodes ordered 
                Reply  Fri Mar 17 15:30:39 2017, Steve, Summary, SUS, oplev laser summary updated oplev_sums.png
Message ID: 12341     Entry time: Wed Jul 27 11:40:48 2016     In reply to: 12126     Reply to this: 12753   12892
Author: steve 
Type: Summary 
Category: SUS 
Subject: oplev laser summary updated 

 

Quote:

 

Quote:

 

Quote:

 

Quote:

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

      May  13, 2014             ETMX,  .............laser in place 90 d

      May  22, 2012             ETMY, 

     Oct.  7,  2013             ETMY,  LT  503 d  or  1.4 y............bad beam quality ?

     Aug. 8,  2014              ETMY,  .............laser in place   425 days  or  1.2 y

 

      Sept. 5, 2014              new 1103P, sn P893516  installed at SP table for aLIGO oplev use qualification

     

           May 23, 2016             ITMX dead laser sn P845648 replaced after 1062 days [2.9 yrs] by 1103P, sn P859884, with output output  2.6 mW, nicely round beam quality at 15 meters.

               July 27, 2016             2  new 1103P from Edmonds in: P947034 & P947039, manf. date April 2016,

ELOG V3.1.3-